OpenStudy (anonymous):

Solve 3x + 2y = –12 for y.

OpenStudy (anonymous):

3x+2y=-12 2y=-12-3x y=(-12-3x)/2

OpenStudy (anonymous):

y=-3

OpenStudy (uri):

3x+2y=-12 2y =-12/3x 2y=-4x y=-4x/2 y=-2x